I really need your help in solving this problem. I appreciate your time and efforts. Have a terrific day :) Here is the three part question?
One sample has a mean of M = 6 and a second sample has a mean of M = 12. The two samples are combined into a single set of scores.
A. What is the mean for the combined set if both of the original samples have n = 5 scores?
B. What is the mean for the combined set if the first sample has n = 4 scores and the second sample has n = 8?
C. What is the mean for the combined set if the first sample has n = 8 scores and the second sample has n = 4?

Answer by macston(5194)

Part A: First sample with M=6 and 5 scores Second sample M=12 with 5 scores
With average (mean)=6 5 scores would sum (5 x 6)=30
With average (mean)=12 5 scores would sum (5 x 12)=60
total points=30+60=90 total scores (5+5)=10 scores
so combined mean is 90total points/10 scores=combined M=9 ANSWER A Mean=9
Part B:
With average (mean)=6 4 scores would sum (4 x 6)=24
With average (mean)=12 8 scores would sum (8 x 12)=96
Total points=24+96=120; total scores (4+8)=12 scores
so combined mean is 120 total points/12 scores=combined M=10 ANSWER B Mean=10
Part C:
With average (mean)=6 8 scores would sum (8 x 6)=48
With average (mean)=12 4 scores would sum (4 x 12)=48
Total points=48+48=96; total scores (8+4)=12 scores
so combined mean is 96 total points/12 scores=combined M=8 ANSWER C Mean=8